# Stable Gradients for Stable Learning at Scale in Deep Reinforcement Learning

### v1

@@ 5 Beyond the ALE and PQN
To evaluate the generality of our findings, we extend our analyses. Specifically, we: (i) apply our proposed methods to PPO (Schulman et al., 2017) on the full ALE and on continuous control tasks in Isaac Gym (Makoviychuk et al., 2021); (ii) assess the impact of richer convolutional encoders by replacing the standard CNN backbone used in the ALE with the Impala CNN architecture (Espeholt et al., 2018); and (iii) augment Simba (Lee et al., 2025) with our proposed techniques and evaluate performance on the DeepMind Control Suite (DMC) (Tassa et al., 2018).
@@
Figure 8 (left) shows that augmenting PPO with the same strategies as in PQN (Layer Normalization by default on PQN, multi-skip residual connections, and Kronecker-factored optimization) significantly boots performance. On the ALE benchmark, the augmented PPO outperforms the baseline in 83.64% of the environments, achieving a median relative improvement of 31.40%. In Isaac Gym’s continuous control tasks, including Cartpole and Anymal (Figure 8, right), the baseline PPO collapses as model size increases, while the augmented variant remains stable and achieves superior performance at all depths and widths.
@@
The Impala CNN is a scalable convolutional architecture that has demonstrated strong performance gains in agents such as Impala (Espeholt et al., 2018) and Rainbow (Hessel et al., 2018). We investigate whether, given its capacity to extract richer representations from visual input, combining Impala CNN with our gradient flow interventions enables effective scaling of the MLP component. As shown in Figure 9, PPO and PQN benefit significantly from replacing the standard CNN with the Impala CNN. For PQN, the Impala encoder enables successful scaling of the MLP, in contrast to the performance collapse seen without our interventions. These results suggest that the expressivity of richer visual encoders is more effectively leveraged by deeper networks when gradient flow is preserved.
@@
Simba (Lee et al., 2025) is a scalable actor-critic framework that integrates normalization, residual connections, and LayerNorm. We augment Simba by replacing its default AdamW optimizer with Kron while keeping all other hyperparameters fixed. We evaluate SAC (Haarnoja et al., 2018) and DDPG (Lillicrap et al., 2015) on challenging DMC tasks, using the Simba architectures of varying depth and width. Despite its design for scalability, default Simba collapses across all tasks as networks grow as shown in Figure 10 (additional results in subsection C.5). In contrast, the Kron-augmented version successfully scales in both depth and width, achieving consistent and stable performance gains. These findings underscore the generality of our approach as effectively enabling parameter scaling in deep RL agents.
Figure 10: Performance comparison between AdamW (dashed lines) and Kron (solid lines) optimizers using the SimBa architecture with SAC and DDPG, averaged over 5 random seeds. As model size increases, AdamW leads to consistent performance degradation, while Kron enables stable and improved learning with larger networks.
@@ 6 Related Work
A central challenge in scaling deep RL lies in the inefficient use of model capacity. Increasing parameter counts often fails to yield proportional gains due to under-utilization. Sokar et al. (2023) show that online RL induces a growing fraction of inactive neurons, a phenomenon also observed in offline settings. Ceron et al. (2024a) report that up to 95% of parameters can be pruned post-training with negligible performance drop, underscoring substantial redundancy. These findings have motivated techniques such as weight resetting (Schwarzer et al., 2023), tokenized computation (Sokar et al., 2025), and sparse architectures (Ceron et al., 2024b, Liu et al., 2025, Willi et al., 2024), along with auxiliary objectives to promote capacity utilization (Farebrother et al., 2023). While scaling model size offers greater expressivity, its benefits depend on appropriate training strategies (Ota et al., 2021). Architectural interventions such as SimBa (Lee et al., 2025) improve robustness by regularizing signal propagation through components such as observation normalization, residual feedforward blocks, and layer normalization. Complementarily, BRO (Nauman et al., 2024) shows that scaling the critic network yields substantial gains in sample and compute efficiency, provided it is paired with strong regularization and optimistic exploration strategies.
Gradient flow, however, remains a central bottleneck. We complement prior efforts by explicitly targeting vanishing gradients as a mechanism for improving scalability. Our approach builds on the role of LayerNorm in stabilizing training and enhancing plasticity (Lyle et al., 2024), and leverages its theoretical effect on gradient preservation as formalized in PQN (Gallici et al., 2025). Optimization-level interventions such as second-order methods (Martens and Grosse, 2015, Muppidi et al., 2024) and adaptive optimizers (Ellis et al., 2024, Bengio et al., 2021, Wu et al., 2017) also address instability under non-stationarity. Our approach integrates architectural and optimizer-level interventions to enable stable gradient flow and unlock parameter scaling in deep RL agents.
@@ 7 Discussion
Our analyses in section 3 suggest that the difficulty in scaling networks in deep RL stems from the interaction between inherent non-stationarity and gradient pathologies that worsen with network size. In section 4, we introduced targeted interventions to address these challenges, and in subsection 4.3, we demonstrated their effectiveness. We validated the generality of our approach across agents and environment suites, consistently observing similar trends. These findings reaffirm the critical role of network design and optimization dynamics in training scalable RL agents. While our proposed solutions may not be optimal, they establish a strong baseline and provide a foundation for future work on gradient stabilization in deep RL. More broadly, our findings suggest that scaling limitations in deep RL are not solely attributable to algorithmic instability or insufficient exploration, but also stem from gradient pathologies amplified by architectural and optimization choices. Addressing these issues directly, without altering the learning algorithm, yields substantial gains in scalability and performance. This suggests that ensuring stable gradient flow is a necessary precondition for effective parameter scaling in deep RL.
Limitations. Our study is constrained by computational resources, which limited our ability to explore architectures beyond a certain size. While our interventions show consistent improvements across agents and environments, further scaling remains an open question. While using second order optimizers introduced additional computational overhead (see Table 10), this cost is mitigated by leveraging vectorized environments and efficient deep RL algorithms, narrowing the gap relative to standard methods. These limitations highlight promising directions for future work, including the development of more computationally efficient gradient stabilization strategies and scalable optimization techniques.
8 Acknowledgment
The authors would like to thank João Guilherme Madeira Araújo, Evan Walters, Olya Mastikhina, Dhruv Sreenivas, Ali Saheb Pasand, Ayoub Echchahed and Gandharv Patil for valuable discussions during the preparation of this work. João Araújo deserves a special mention for providing us valuable feed-back on an early draft of the paper. We want to acknowledge funding support from Google, CIFAR AI and compute support from Digital Research Alliance of Canada and Mila IDT. We would also like to thank the Python community (Van Rossum and Drake Jr, 1995, Oliphant, 2007) for developing tools that enabled this work, including NumPy Harris et al. (2020), Matplotlib (Hunter, 2007), Jupyter (Kluyver et al., 2016), and Pandas (McKinney, 2013).
@@ Appendix A Environment Details
Throughout the paper, we evaluate the deep reinforcement learning agents’ performance on the Atari-10 suite (Aitchison et al., 2023), a curated subset of games from the Arcade Learning Environment (ALE) (Bellemare et al., 2013). Atari-10 consists of 10 games selected to capture the maximum variance in algorithm performance, achieving over 90% correlation with results on the full ALE benchmark. This makes it a computationally efficient yet representative testbed for deep reinforcement learning. We follow the experimental protocol of Obando Ceron et al. (2023), Ceron et al. (2024b), Agarwal et al. (2021), running each experiment with three random seeds and reporting the aggregate human-normalized score across games.
@@
Additionally, to further support the generality of our findings, we evaluate the proposed combined gradient interventions on the full ALE benchmark. We also assess their effectiveness on continuous control tasks from the IsaacGym simulator (Makoviychuk et al., 2021) and the DeepMind Control Suite (DMC) (Tassa et al., 2018), extending our analysis to robotics-based environments. We conduct experiments on the 4 challenging tasks of DMC:
@@ Appendix B Network Sizes
Throughout the paper, we experiment with models of varying depths and widths. Unless stated otherwise (e.g. in  section 5, where we evaluate the Impala CNN), the convolutional feature extractors are kept fixed. Consequently, our experiments focus primarily on scaling strategies and architectural variations in the MLP components of the networks.
@@
Depth / Width Small Medium Large
Small 2.39 11.90 27.70
Medium 3.45 21.35 53.93
Large 4.50 30.79 80.15
@@ C.1 Scaling with DQN and Rainbow
To further support our hypothesis on the emergence of gradient pathologies in deep reinforcement learning, we investigate whether similar issues arise in algorithms beyond PQN and PPO, as discussed in the main paper. Specifically, we study the effects of architectural scaling on two widely used value-based algorithms: DQN (Mnih et al., 2015) and Rainbow (Hessel et al., 2018).
@@
In Figure 11, we report the performance of DQN and Rainbow as we scale the depth and width of their networks. As with PQN and PPO, we observe consistent degradation in performance at larger scales. In Figure 12, we present the corresponding gradient behavior, which reveals the same vanishing and destabilization phenomena discussed in this work. These findings reinforce the generality of the identified gradient pathologies across both policy-based and value-based deep RL algorithms.
@@ C.2 Combining Gradient Interventions in Non-stationary Supervised Learning
Building on our findings in subsection 4.3, we extend our analysis by applying the proposed combined gradient interventions to the same image classification models used in section 3. Specifically, we train the models in the non-stationary supervised learning setup, where the CIFAR-10 labels are iteratively shuffled, following the experimental design from Sokar et al. (2023). As demonstrated in section 3, while models in standard supervised learning settings are able to scale effectively and maintain high performance, introducing non-stationarity leads to failure in adaptation for baselines that use fully connected layers and the Adam optimizer. This issue is exacerbated as model scale increases.
Our results, presented in Figure 7, show that combining the multi-skip architecture for the MLP component with the Kronecker-factored optimizer and Layer Normalization enables near-perfect continuous adaptation. The models quickly adapt to the changing optimization problem following label reshuffling, with gradient magnitudes remaining stable throughout the process.
@@ C.3 Architecture and Optimizer Ablations
In this work, we introduce the multi-skip architecture, an extension of the standard residual MLP design, and propose the use of the Kronecker-factored optimizer for online deep RL. While these techniques form the basis of our primary interventions, our broader goal is not to prescribe a fixed set of methods, but rather to motivate a general class of architectural and optimization interventions that promote healthy gradient flow in deep networks. To this end, we expand the scope of our evaluation by incorporating a wider range of baselines. Specifically, we compare various optimizer choices, including Adam and AdaBelief (Zhuang et al., 2020), alongside MLP architectures such as the standard residuapl MLP (He et al., 2016) and DenseNet (Huang et al., 2017). These architectures have been previously explored in the context of scaling networks in online deep RL (Lee et al., 2025, Ota et al., 2024), providing a relevant basis for comparison.
We also evaluated state-of-the-art optimizers that have demonstrated success in training large-scale models such as transformers in supervised learning. Specifically, we tested Shampoo (Gupta et al., 2018), a second-order optimizer that maintains and preconditions gradients using full-matrix statistics per layer, and Apollo (Ma, 2020), an adaptive optimizer that leverages curvature information without explicitly computing or storing second-order matrices.
Despite extensive hyperparameter tuning for both methods, we were unable to achieve strong performance in the online deep RL setting. This suggests that further investigation is needed to understand the key properties required for these optimizers to be effective in this regime. Asadi et al. (2023), Ceron and Castro (2021) demonstrate that optimizer behavior plays a critical role in the training dynamics of online deep RL methods, with Asadi et al. (2023) showing that stale optimizer states can hinder learning, and Ceron and Castro (2021) revealing that optimizer sensitivity interacts with the choice of loss function, particularly when comparing Huber and MSE losses.
We present the results for PPO and PQN across all tested optimizers in Figure 13.
@@
Results with the Multi-Skip Architecture.
We present the full learning curves comparing the proposed multi-skip architecture to the baseline fully connected architecture across all depths and widths studied in the paper. We follow the experimental protocol of Obando Ceron et al. (2023), Ceron et al. (2024b), Agarwal et al. (2021), running each experiment with three random seeds.
@@
Results with the Kron Optimizer.
We present the full learning curves comparing the Kron optimizer to the baseline RAdam optimizer originally used in PQN (Gallici et al., 2025), across all depths and widths studied in the paper. We follow the experimental protocol of Obando Ceron et al. (2023), Ceron et al. (2024b), Agarwal et al. (2021), running each experiment with three random seeds.
@@
C.4 Results on the Full ALE
In this section, we provide the full training curves corresponding to the aggregated results shown in subsection 4.3, where we evaluate the performance of the PQN and PPO agents on the full set of environments from the ALE after applying our two proposed gradient interventions. The per-environment learning curves are presented in Figure 16 for PQN and Figure 17 for PPO. We follow the experimental protocol of Obando Ceron et al. (2023), Ceron et al. (2024b), Agarwal et al. (2021), running each experiment with three random seeds.
C.5 Simba on DMC
In this section, we present the full results accompanying the experiments combining Simba (Lee et al., 2025) with our proposed gradient interventions, as introduced in subsection 4.3. For these experiments, we retain Simba’s original architectural choices but replace the AdamW optimizer with Kron.
The following figures illustrate these findings:
Figure 18: SAC scaling depth
Figure 19: SAC scaling width
Figure 20: SAC scaling both depth and width
Figure 21: DDPG scaling depth
Figure 22: DDPG scaling width
Figure 23: DDPG scaling both depth and width

### v2

@@ 5 Beyond the ALE and PQN
To evaluate the generality of our findings, we extend our analyses. Specifically, we: (i) apply our proposed methods to PPO (Schulman et al., 2017) on the full ALE and on continuous control tasks in Isaac Gym (Makoviychuk et al., 2021); (ii) assess the impact of richer convolutional encoders by replacing the standard CNN backbone used in the ALE with the Impala CNN architecture (Espeholt et al., 2018); (iii) augment Simba (Lee et al., 2025) with our proposed techniques and evaluate performance on the DeepMind Control Suite (DMC) (Tassa et al., 2018); and (iv) investigate whether our interventions can stabilize and scale a wide range of Q-learning algorithms in challenging offline RL benchmarks (Park et al., 2025).
@@
Fig. 8 (left) shows that augmenting PPO with the same strategies as in PQN (Layer Normalization by default on PQN, multi-skip residual connections, and Kronecker-factored optimization) significantly boots performance. On the ALE benchmark, the augmented PPO outperforms the baseline in 83.64% of the environments, achieving a median relative improvement of 31.40%. In Isaac Gym’s continuous control tasks, including Cartpole and Anymal (Fig. 8, right), the baseline PPO collapses as model size increases, while the augmented variant remains stable and achieves superior performance at all depths and widths.
@@
The Impala CNN is a scalable convolutional architecture that has demonstrated strong performance gains in agents such as Impala (Espeholt et al., 2018) and Rainbow (Hessel et al., 2018). We investigate whether, given its capacity to extract richer representations from visual input, combining Impala CNN with our gradient flow interventions enables effective scaling of the MLP component. As shown in Fig. 9, PPO and PQN benefit significantly from replacing the standard CNN with the Impala CNN. For PQN, the Impala encoder enables successful scaling of the MLP, in contrast to the performance collapse seen without our interventions. These results suggest that the expressivity of richer visual encoders is more effectively leveraged by deeper networks when gradient flow is preserved.
@@
Simba (Lee et al., 2025) is a scalable actor-critic framework that integrates normalization, residual connections, and LayerNorm. We augment Simba by replacing its default AdamW optimizer with Kron while keeping all other hyperparameters fixed. We evaluate SAC (Haarnoja et al., 2018) and DDPG (Lillicrap et al., 2015) on challenging DMC tasks, using the Simba architectures of varying depth and width. Despite its design for scalability, default Simba collapses across all tasks as networks grow as shown in Fig. 10 (additional results in Sec. E.2). In contrast, the Kron-augmented version successfully scales in both depth and width, achieving consistent and stable performance gains. These findings underscore the generality of our approach as effectively enabling parameter scaling in deep RL agents.
Figure 10: Performance comparison between AdamW (dashed lines) and Kron (solid lines) optimizers using the SimBa architecture with SAC and DDPG, averaged over 55 random seeds. As model size increases, AdamW leads to consistent performance degradation, while Kron enables stable and improved learning with larger networks.
Gradient Interventions for Scalable Offline Q-Learning.
Park et al. (2025) highlight significant challenges in scaling Q-learning algorithms for complex offline RL tasks, demonstrating that many standard offline RL baselines fail to learn effective policies, even on large, high-quality datasets. Their key finding was that performance improvements were primarily driven by techniques that shorten the effective credit assignment horizon, such as n-step returns and hierarchical methods. This led to their proposal of two new high-performing algorithms, SHARSA and DSHARSA, which are designed to operate with shorter effective horizons. This finding motivates a parallel investigation: can our proposed gradient interventions, which are designed to stabilize and accelerate deep network training, also address the scaling limitations of offline Q-learning? To test this, we augment the full suite of baselines and novel algorithms from Park et al. (2025) with our proposed gradient interventions. The results, presented in Fig. 11, show that our methods provide a complementary path to scalability. Applying our interventions generally improves the performance of the baselines across all tasks.
The performance gains are particularly pronounced in the most sparse-reward task, humanoidmaze-giant-navigate, where our gradient interventions enable multiple methods to achieve near-optimal performance, whereas their baseline counterparts largely fail. Furthermore, this stability extends to generalization across task difficulty. When moving from puzzle-4x5-play to the harder puzzle-4x6-play task, many baselines exhibit a sharp performance degradation. In contrast, the performance of several algorithms with our interventions remains consistent and high, demonstrating improved robustness. Finally, we note that while the primary focus of this paper is to address gradient pathologies arising from scaling and non-stationarity, these results highlight that our interventions are also highly beneficial in offline deep RL, where inputs are stationary.
Figure 11: Performance of offline Q-learning algorithms with and without our gradient interventions. We compare the original algorithms from Park et al. (2025) against our augmented versions. The results, averaged over 33 seeds demonstrate a general improvement in scalability.
@@ 6 Related Work
A central challenge in scaling deep RL lies in the inefficient use of model capacity. Increasing parameter counts often fails to yield proportional gains due to under-utilization. Sokar et al. (2023); Mayor et al. (2025) and Liu et al. (2025c) show that online RL induces a growing fraction of inactive neurons, a phenomenon also observed in offline settings. Ceron et al. (2024b) report that up to 95% of parameters can be pruned post-training with negligible performance drop, underscoring substantial redundancy. These findings have motivated techniques such as weight resetting (Schwarzer et al., 2023), tokenized computation (Sokar et al., 2025), and sparse architectures (Ceron et al., 2024c; Willi et al., 2024; Liu et al., 2025a; Ma et al., 2025), along with auxiliary objectives to promote capacity utilization (Farebrother et al., 2023). While scaling model size offers greater expressivity, its benefits depend on appropriate training strategies (Ota et al., 2021). Architectural interventions such as SimBa (Lee et al., 2025) improve robustness by regularizing signal propagation through components such as observation normalization, residual feedforward blocks, and layer normalization. Complementarily, BRO (Nauman et al., 2024) shows that scaling the critic network yields substantial gains in sample and compute efficiency, provided it is paired with strong regularization and optimistic exploration strategies.
Gradient flow, however, remains a central bottleneck. We complement prior efforts by explicitly targeting vanishing gradients as a mechanism for improving scalability. Our approach builds on the role of LayerNorm in stabilizing training and enhancing plasticity (Lyle et al., 2024), and leverages its theoretical effect on gradient preservation as formalized in PQN (Gallici et al., 2025). Optimization-level interventions such as second-order methods (Martens and Grosse, 2015; Muppidi et al., 2024) and adaptive optimizers (Ellis et al., 2024; Bengio et al., 2021; Wu et al., 2017) also address instability under non-stationarity. Our approach integrates architectural and optimizer-level interventions to enable stable gradient flow and unlock parameter scaling in deep RL agents.
@@ 7 Discussion
Our analyses in Sec. 3 suggest that the difficulty in scaling networks in deep RL stems from the interaction between inherent non-stationarity and gradient pathologies that worsen with network size. In Sec. 4, we introduced targeted interventions to address these challenges, and in Sec. 4.3, we demonstrated their effectiveness. We validated the generality of our approach across agents and environment suites, consistently observing similar trends. These findings reaffirm the critical role of network design and optimization dynamics in training scalable RL agents. While our proposed solutions may not be optimal, they establish a strong baseline and provide a foundation for future work on gradient stabilization in deep RL. More broadly, our findings suggest that scaling limitations in deep RL are not solely attributable to algorithmic instability or insufficient exploration, but also stem from gradient pathologies amplified by architectural and optimization choices. Addressing these issues directly, without altering the learning algorithm, yields substantial gains in scalability and performance. This suggests that ensuring stable gradient flow is a necessary precondition for effective parameter scaling in deep RL.
Limitations. Our study is constrained by computational resources, which limited our ability to explore architectures beyond a certain size. While our interventions show consistent improvements across agents and environments, further scaling remains an open question. While using second order optimizers introduced additional computational overhead (see Tab. 12), this cost is mitigated by leveraging vectorized environments and efficient deep RL algorithms, narrowing the gap relative to standard methods. These limitations highlight promising directions for future work, including the development of more computationally efficient gradient stabilization strategies and scalable optimization techniques.
Acknowledgment
The authors would like to thank João Guilherme Madeira Araújo, Evan Walters, Olya Mastikhina, Dhruv Sreenivas, Ali Saheb Pasand, Ayoub Echchahed and Gandharv Patil for valuable discussions during the preparation of this work. João Araújo deserves a special mention for providing us valuable feed-back on an early draft of the paper. We want to acknowledge funding support from Natural Sciences and Engineering Research Council (NSERC) of Canada, Google Research, Fonds de recherche du Québec (FRQNT) and The Canadian Institute for Advanced Research (CIFAR) and compute support from Digital Research Alliance of Canada, Mila IDT, and NVidia. We would also like to thank the Python community Van Rossum and Drake Jr (1995); Oliphant (2007) for developing tools that enabled this work, including NumPy Harris et al. (2020), Matplotlib Hunter (2007), Jupyter Kluyver et al. (2016), and Pandas McKinney (2013).
Broader Impact
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Environment Details
Throughout the paper, we evaluate the deep reinforcement learning agents’ performance on the Atari-10 suite [Aitchison et al., 2023], a curated subset of games from the Arcade Learning Environment (ALE) [Bellemare et al., 2013]. Atari-10 consists of 10 games selected to capture the maximum variance in algorithm performance, achieving over 90% correlation with results on the full ALE benchmark. This makes it a computationally efficient yet representative testbed for deep reinforcement learning. We follow the experimental protocol of Obando Ceron et al. [2023], Ceron et al. [2024c], Agarwal et al. [2021], running each experiment with three random seeds and reporting the aggregate human-normalized score across games.
@@
Additionally, to further support the generality of our findings, we evaluate the proposed combined gradient interventions on the full ALE benchmark. We also assess their effectiveness on continuous control tasks from the IsaacGym simulator [Makoviychuk et al., 2021] and the DeepMind Control Suite (DMC) [Tassa et al., 2018], extending our analysis to robotics-based environments. We conduct experiments on the 4 challenging tasks of DMC:
@@ Appendix B Network Sizes
Throughout the paper, we experiment with models of varying depths and widths. Unless stated otherwise (e.g. in  Sec. 5, where we evaluate the Impala CNN), the convolutional feature extractors are kept fixed. Consequently, our experiments focus primarily on scaling strategies and architectural variations in the MLP components of the networks.
@@
[…]
@@ C.1 Scaling with DQN and Rainbow
To further support our hypothesis on the emergence of gradient pathologies in deep reinforcement learning, we investigate whether similar issues arise in algorithms beyond PQN and PPO, as discussed in the main paper. Specifically, we study the effects of architectural scaling on two widely used value-based algorithms: DQN [Mnih et al., 2015] and Rainbow [Hessel et al., 2018].
@@
In Fig. 12, we report the performance of DQN and Rainbow as we scale the depth and width of their networks. As with PQN and PPO, we observe consistent degradation in performance at larger scales. In Fig. 13, we present the corresponding gradient behavior, which reveals the same vanishing and destabilization phenomena discussed in this work. These findings reinforce the generality of the identified gradient pathologies across both policy-based and value-based deep RL algorithms.
(a) DQN HNS
(b) Rainbow HNS
@@ C.2 Combining Gradient Interventions in Non-stationary Supervised Learning
Building on our findings in Sec. 4.3, we extend our analysis by applying the proposed combined gradient interventions to the same image classification models used in Sec. 3. Specifically, we train the models in the non-stationary supervised learning setup, where the CIFAR-10 labels are iteratively shuffled, following the experimental design from Sokar et al. [2023]. As demonstrated in Sec. 3, while models in standard supervised learning settings are able to scale effectively and maintain high performance, introducing non-stationarity leads to failure in adaptation for baselines that use fully connected layers and the Adam optimizer. This issue is exacerbated as model scale increases.
Our results, presented in Fig. 7, show that combining the multi-skip architecture for the MLP component with the Kronecker-factored optimizer and Layer Normalization enables near-perfect continuous adaptation. The models quickly adapt to the changing optimization problem following label reshuffling, with gradient magnitudes remaining stable throughout the process.
@@ C.3 Architecture and Optimizer Ablations
In this work, we introduce the multi-skip architecture, an extension of the standard residual MLP design, and propose the use of the Kronecker-factored optimizer for online deep RL. While these techniques form the basis of our primary interventions, our broader goal is not to prescribe a fixed set of methods, but rather to motivate a general class of architectural and optimization interventions that promote healthy gradient flow in deep networks. To this end, we expand the scope of our evaluation by incorporating a wider range of baselines. Specifically, we compare various optimizer choices, including Adam and AdaBelief [Zhuang et al., 2020], alongside MLP architectures such as the standard residuapl MLP [He et al., 2016] and DenseNet [Huang et al., 2017]. These architectures have been previously explored in the context of scaling networks in online deep RL [Lee et al., 2025, Ota et al., 2024], providing a relevant basis for comparison.
We also evaluated state-of-the-art optimizers that have demonstrated success in training large-scale models such as transformers in supervised learning. Specifically, we tested Shampoo [Gupta et al., 2018], a second-order optimizer that maintains and preconditions gradients using full-matrix statistics per layer, and Apollo [Ma, 2020], an adaptive optimizer that leverages curvature information without explicitly computing or storing second-order matrices.
Despite extensive hyperparameter tuning for both methods, we were unable to achieve strong performance in the online deep RL setting. This suggests that further investigation is needed to understand the key properties required for these optimizers to be effective in this regime. Asadi et al. [2023], Ceron and Castro [2021] demonstrate that optimizer behavior plays a critical role in the training dynamics of online deep RL methods, with Asadi et al. [2023] showing that stale optimizer states can hinder learning, and Ceron and Castro [2021] revealing that optimizer sensitivity interacts with the choice of loss function, particularly when comparing Huber and MSE losses.
We present the results for PPO and PQN across all tested optimizers in Fig. 14.
@@
C.4 Results with the Multi-Skip Architecture.
We present the full learning curves comparing the proposed multi-skip architecture to the baseline fully connected architecture across all depths and widths studied in the paper. We follow the experimental protocol of Obando Ceron et al. [2023], Ceron et al. [2024c], Agarwal et al. [2021], running each experiment with three random seeds.
@@
C.5 Ablation on the number of skip connections.
To isolate the effect of skip length on performance, we fix the main network of our proposed MultiSkip architecture (in large size, which includes 5 residual blocks) and vary how many of these blocks receive skip connections from the encoder. When Skip = $k$, we apply the encoder features as skip connections to the first $k$ residual blocks immediately following the encoder, while the remaining (5k)(5-k) blocks operate without direct encoder input. The table below reports human-normalized scores on the Atari-10 benchmark Agarwal et al. [2021].
Table 2: Human-normalized scores on the Atari-10 benchmark, varying the number of residual blocks ($k$) that receive skip connections from the encoder. Performance generally improves as more connections are added.
Environment Skip=1 Skip=2 Skip=3 Skip=4 Skip=5
Amidar-v5 0.20 0.17 0.19 0.20 0.36
BattleZone-v5 0.01 0.67 0.62 0.60 0.69
Bowling-v5 0.07 0.04 0.04 0.08 0.23
DoubleDunk-v5 -2.09 -2.00 -1.55 -1.36 -1.32
Frostbite-v5 0.67 0.70 0.79 0.92 0.88
KungFuMaster-v5 0.93 0.95 0.93 1.12 0.98
NameThisGame-v5 0.79 0.65 0.72 0.85 1.24
Phoenix-v5 0.69 0.68 0.68 0.66 0.66
Qbert-v5 0.84 1.01 1.06 0.91 1.09
Riverraid-v5 0.42 0.44 0.65 0.69 0.99
Aggregate (mean) 0.25 0.33 0.41 0.47 0.58
Performance steadily improves as more skip connections are added, peaking when all 5 blocks are connected. This supports our original design decision to broadcast features to all MultiSkip blocks.
C.6 Results with the Kron Optimizer.
We present the full learning curves comparing the Kron optimizer to the baseline RAdam optimizer originally used in PQN [Gallici et al., 2025], across all depths and widths studied in the paper. We follow the experimental protocol of Obando Ceron et al. [2023], Ceron et al. [2024c], Agarwal et al. [2021], running each experiment with three random seeds.
@@
C.7 Justification for Selected Interventions.
The choice of Kronecker-factored optimization and the MultiSkip architecture was the result of a systematic exploration of candidate interventions aimed at improving gradient flow, plasticity, and stability in deep RL. We evaluated a wide range of alternative methods from prior work known to mitigate optimization pathologies [Moalla et al., 2024, Juliani and Ash, 2024].
Our evaluated methods, summarized in Tab. 3, included:
  • Second-order and adaptive optimizers: Apollo [Ma, 2020], Shampoo [Gupta et al., 2018], AdaBelief [Zhuang et al., 2020].
  • Regularization: L2 norm penalties [Kumar et al., 2023], weight clipping, and weight decay [Elsayed et al., 2024].
  • Activation functions: GELU [Hendrycks, 2016] and CReLU [Abbas et al., 2023].
  • Learning rate schedules: Cosine annealing and cyclic schedulers.
  • Learning rate scaling: Multiplying and dividing the default learning rate (2.5e-4) by 10 to compensate for increased network scale.
As shown in the table, none of these interventions consistently improved performance compared to our proposed combination. This motivated our decision to focus on the combination of Kronecker-factored optimization and the multi-skip architecture.
Table 3: Comparison of mean human-normalized scores on Atari-10 for alternative interventions. Our proposed method (Ours) is compared against ablations using only a single alternative intervention. The results (3 seeds per experiment) show that no single alternative consistently matches the performance of our combined approach.
Environment Ours Cosine LR Cyclic LR GELU CReLU L2 Norm Weight Clip Weight Decay LR=2.08e-5 LR=3.00e-3
Amidar-v5 0.355897 -0.001663 0.002276 0.015143 0.016514 0.029293 0.012283 0.013013 0.018848 0.078660
BattleZone-v5 0.694566 0.031297 -0.000287 0.015505 -0.004594 -0.004594 0.035604 0.061446 0.019812 0.575407
Bowling-v5 0.231831 0.050145 -0.008721 0.028706 -0.009811 -0.000727 0.030523 0.051235 -0.086483 0.067587
DoubleDunk-v5 -1.318182 -2.318182 -2.318182 -2.363636 -2.318182 -2.363636 -2.454545 -2.454545 -2.454545 -2.090909
Frostbite-v5 0.881087 -0.001921 0.027708 0.008890 -0.010938 0.004755 0.075372 0.021736 0.008034 0.110036
KungFuMaster-v5 0.975251 -0.011055 0.003181 -0.000823 -0.011500 -0.011278 -0.011500 -0.011055 -0.011055 1.362077
NameThisGame-v5 1.242066 -0.030191 -0.132767 -0.261400 -0.198430 -0.068147 -0.197995 -0.148227 -0.134417 0.747251
Phoenix-v5 0.655141 0.072687 -0.087237 -0.090014 -0.090168 -0.068722 -0.090940 -0.093949 -0.102358 0.411968
Qbert-v5 1.094894 0.013814 0.010052 -0.002833 0.008359 0.006102 0.016259 0.000177 0.001305 0.112282
Riverraid-v5 0.993568 -0.000127 0.345100 0.304636 -0.056846 -0.058494 0.293514 -0.044456 0.057765 0.237365